OpenStudy (anonymous):

How will I solve for x and y, Given that: 4x-2=8 and y=(x-8)

OpenStudy (anonymous):

Solve 4x-2 = 8 first by adding 2 to be sides to isolate x So, it become 4x = 10 x = 2.5 Plug that value in y = (x-8) y = (2.5-8) = -5.5

OpenStudy (anonymous):

*both sides